# Prognostic Framework for Robotic Manipulators Operating Under Dynamic Task Severities

### v2

@@ II-B Prognostics Under Dynamic Operating Conditions
One of the key research gaps identified in the reviewed literature was the lack of a true closed-form approximation for the residual lifetime distribution (RLD) under dynamic operating conditions. Existing works such as [22] and [23] employed Laplace transform techniques to derive analytical expressions for the first-passage time of a Brownian motion process in a Markovian environment. However, these expressions exist only in the Laplace domain and require numerical inversion (e.g., the Abate–Whitt algorithm), which limits numerical stability and scalability for real-time prediction. More importantly, these formulations assume that both the degradation process and the environmental state are continuously observable. In contrast, our framework considers a partially observed setting in which the degradation is measured only at discrete inspection epochs, while the task-severity environment evolves continuously during the operational cycles between inspections. Since a robot may remain idle for extended periods, degradation (position accuracy) observations are collected only at inspection epochs after a fixed number of operational cycles. The data collection procedure is discussed in Section III.
@@ IV-A Approach 1: Effective Degradation Rate
Note that in this approach, all information from the CTMC generator $Q$ enters the RLD only through its stationary distribution $\pi$. The effective degradation rate $\mu_{\phi}$ depends on the $Q$ matrix solely via $\pi$, which satisfies $\pi Q=0$. Consequently, any mismatch between the true generator $Q^{\star}$ and an estimated generator $\widehat{Q}$ affects the predicted RLD if and only if it alters the corresponding stationary distribution. If $\widehat{Q}$ and $Q^{\star}$ yield the same $\pi$, then the RLD remains invariant to the $Q$-mismatch. Appendix B provides the mathematical relation that expresses the mismatch in $Q$ as a function of the induced mismatch in $\pi$.
@@ IV-D Theoretical Analysis
Please refer to the Appendix C for proof. ∎
@@ V-C Case Study 1: Planar Robotic Simulator
The parameterization of the model follows the published specifications of the UR5e robot (Table IV), with all degrees of freedom locked except the joints of interest. The controller consists of a gravity compensator and a linear-quadratic regulator (LQR). The LQR method produces an optimal state-feedback controller given a cost function specifying the relative importance of state deviation and minimizing effort (torque). The resulting controller acts at a defined frequency to minimize the difference between the observed and desired configuration of the robot. The costs of the minimization function are also tuned to achieve a reasonable response time with torques well below the stated maximums for the UR5e robot. Given the robot equations of motion with parameters from Table IV, gravity compensator, and LQR controller which can be shown to be stable (see Appendix D), the system was integrated using the LSODA [32] method from SciPy. This method switches between stiff and nonstiff integration methods which are particularly useful for systems in which small changes can induce large variability or instabilities such as the nonlinear case being explored. We use a 0.005-second integration time step and a 10-second cycle time for each motion. A schematic of the planar robot is given in Figure 3. For the planar robot, a task is defined as a complete pick-and-place motion between two fixed joint configurations. Let the joint angles be defined as $\theta_{1}$ and $\theta_{2}$, which correspond to the upper arm and forearm, respectively. In each task, the robot is commanded to move from the joint configuration of $(\theta_{1},\theta_{2})=(10^{\circ},90^{\circ})$ to $(\theta_{1},\theta_{2})=(90^{\circ},20^{\circ})$.
@@ V-D Case Study 2: Spatial Robotic Simulator
An operational space controller is used, which commands torques to move the end-effector to the desired position with low kinematic energy. An end-effector trajectory is linearly interpolated and discretized from the goal pose to the current pose, and a desired wrench of the end-effector is computed at each point. A wrench is a “spatial force” or moment and force expressed in a reference frame. The wrench is then propagated to the joints, given the Jacobian of the current joint configuration. The torque at each joint is clipped to the reported maximum rating, limiting the response of the robot upon reaching high degradation values. For the spatial robot, a task is defined as a complete pick-and-place operation along a prescribed spatial trajectory segment. The robot begins a cycle in a joint configuration with the end effector at the cube with the grips open. Then, a trajectory is assembled which commands end-effector poses at certain times as listed in Table V. This controller has been shown to be stable (see Appendix D). Each cycle is simulated for 5 seconds, enough time for the robot arm to reach the final position.
@@ Appendix A Proof of Proposition 1
Appendix B Effect of $Q$ Matrix Mismatch via $\pi$
For the task severity CTMC with generator $Q\in\mathbb{R}^{S\times S}$, the stationary distribution $\pi$ satisfies
$\pi^{\top}Q=0,\quad\text{and}\quad\pi^{\top}\mathbf{1}=1.$ (35)
Let $Q^{\star}$ denote the true generator and $\widehat{Q}=Q^{\star}+\Delta Q$ an estimated generator, where each row of $\Delta Q$ sums to zero. The corresponding stationary distributions are $\pi^{\star}$ and $\widehat{\pi}=\pi^{\star}+\Delta\pi$, respectively.
Using the group inverse $Q^{\#}$, defined by
$QQ^{\#}=I-\mathbf{1}\pi^{\top},\quad Q^{\#}Q=I-\mathbf{1}\pi^{\top},\quad Q^{\#}\mathbf{1}=0,$ (36)
The perturbation of the stationary distribution satisfies,
$\Delta\pi^{\top}=\pi^{\top}\,\Delta Q\,Q^{\#},$ (37)
with $\Delta\pi^{\top}\mathbf{1}=0$.
Hence, the induced change in the effective degradation rate is
$\Delta\mu_{\phi}=\Delta\pi^{\top}\mu=\pi^{\top}\,\Delta Q\,Q^{\#}\mu.$ (38)
Appendix C Proof of Proposition 2
@@
Appendix D Notes on Stability Analysis

### v3

@@ II-B Prognostics Under Dynamic Operating Conditions
One of the key research gaps identified in the reviewed literature was the lack of closed-form approximation for RLD when subjected to dynamic operating conditions. Addressing this research gap is the primary methodological contribution of this paper. Our methodology follows the formulation developed by [23] with some key differences. First, our robot’s accuracy degradation signal does not contain any shocks. This allows us to propose a closed-form expression of the RLD of the robot that still leverages the properties of the CTMC. Furthermore, our problem setting is unique in that we cannot assume that the robot is constantly being utilized. Thus, the conventional assumption used in most degradation models, i.e., monitoring at fixed time intervals will most likely generate erroneous results. In our problem setting, since a robot can sometimes be idle, we instead collect degradation (position accuracy) observations at the inspection epochs after a fixed number of operational cycles. A discussion on the data collection procedure is given in Section III.
@@ IV-D Theoretical Analysis
Please refer to the Appendix B for proof. ∎
@@ V-C Case Study 1: Planar Robotic Simulator
The parameterization of the model follows the published specifications of the UR5e robot (Table IV), with all degrees of freedom locked except the joints of interest. The controller consists of a gravity compensator and a linear-quadratic regulator (LQR). The LQR method produces an optimal state-feedback controller given a cost function specifying the relative importance of state deviation and minimizing effort (torque). The resulting controller acts at a defined frequency to minimize the difference between the observed and desired configuration of the robot. The costs of the minimization function are also tuned to achieve a reasonable response time with torques well below the stated maximums for the UR5e robot. Given the robot equations of motion with parameters from Table IV, gravity compensator, and LQR controller which can be shown to be stable (see Appendix C), the system was integrated using the LSODA [32] method from SciPy. This method switches between stiff and nonstiff integration methods which are particularly useful for systems in which small changes can induce large variability or instabilities such as the nonlinear case being explored. We use a 0.005-second integration time step and a 10-second cycle time for each motion. A schematic of the planar robot is given in Figure 3. For the planar robot, a task is defined as a complete pick-and-place motion between two fixed joint configurations. Let the joint angles be defined as $\theta_{1}$ and $\theta_{2}$, which correspond to the upper arm and forearm, respectively. In each task, the robot is commanded to move from the joint configuration of $(\theta_{1},\theta_{2})=(10^{\circ},90^{\circ})$ to $(\theta_{1},\theta_{2})=(90^{\circ},20^{\circ})$.
@@ V-D Case Study 2: Spatial Robotic Simulator
An operational space controller is used, which commands torques to move the end-effector to the desired position with low kinematic energy. An end-effector trajectory is linearly interpolated and discretized from the goal pose to the current pose, and a desired wrench of the end-effector is computed at each point. A wrench is a “spatial force” or moment and force expressed in a reference frame. The wrench is then propagated to the joints, given the Jacobian of the current joint configuration. The torque at each joint is clipped to the reported maximum rating, limiting the response of the robot upon reaching high degradation values. For the spatial robot, a task is defined as a complete pick-and-place operation along a prescribed spatial trajectory segment. The robot begins a cycle in a joint configuration with the end effector at the cube with the grips open. Then, a trajectory is assembled which commands end-effector poses at certain times as listed in Table V. This controller has been shown to be stable (see Appendix C). Each cycle is simulated for 5 seconds, enough time for the robot arm to reach the final position.
@@ Appendix A Proof of Proposition 1
Appendix B Proof of Proposition 2
@@
Appendix C Notes on Stability Analysis